OpenStudy (mendicant_bias):

Alright, next problem in the long list that I couldn't figure out how to do using convolution (posted below in a moment).

OpenStudy (mendicant_bias):

\[\mathcal{L^{-1}}\left\{\vphantom{}\frac{1}{s^3(s-1)}\right\}\]

OpenStudy (mendicant_bias):

(Or, excuse me, I'm not supposed to use convolution in the regular sense, but I am supposed to use the F(s)/s convolution property):

OpenStudy (mendicant_bias):

\[\int\limits_{0}^{t}f(\tau)d \tau=\mathcal{L}\left\{\vphantom{}\frac{F(s)}{s}\right\}\](where g(t) or g(t-tau)=1)

OpenStudy (mendicant_bias):

Putting it in this general form:\[\mathcal{L}\left\{\vphantom{}\frac{1/s^2(s-1)}{s}\right\}\]

OpenStudy (mendicant_bias):

\[\mathcal{L^{-1}}\left\{\vphantom{}\frac{1}{s^2(s-1)}\right\}=?\]I can approach this using partial fractions, but I was wondering if somebody had a preferred or better way.

OpenStudy (mendicant_bias):

(Just dealing with F(s) by itself, above)

OpenStudy (mendicant_bias):

@sidsiddhartha , do you think there is a better way to do this inverse laplace at the bottom other than partial fractions, if the 1/s F(s) property has already been applied?

OpenStudy (sidsiddhartha):

\[take\\F_1(s)=\frac{ 1}{ s^3 }\\and\\F_2(s)=\frac{ 1 }{ s-1 }\] then find inverse and then conv.integral

OpenStudy (mendicant_bias):

Oh, so you can do it again? This was one of my earlier questions but I didn't outright ask it; this separating inverse laplace functions and using convolution is recursive, like,

OpenStudy (mendicant_bias):

(Oh well, enough words for now, just going to do it):\[\mathcal{L^{-1}}\left\{F_{1}(s) \right\}=t, \ \ \ \mathcal{L^{-1}}\left\{F_{2}(s)\right\}=e^t\]

OpenStudy (mendicant_bias):

\[F_{1}(s)F_{2}(s)=te^{t}\]

OpenStudy (sidsiddhartha):

hey check again L91/s^3]=?

OpenStudy (sidsiddhartha):

oh typo

OpenStudy (sidsiddhartha):

\[l[1/s^3] \neq t\]

OpenStudy (mendicant_bias):

Oh, read that as a 3, my bad, one sec\[\mathcal{L^{-1}}\left\{\vphantom{}\frac{1}{s^{3}}\right\}=\frac{t^2}{2}\]

OpenStudy (sidsiddhartha):

yeah :)

OpenStudy (mendicant_bias):

Alright, now I'm getting confused because I'm trying to go quick because of time constraint and am prone to make some mistakes, lol, one sec.

OpenStudy (mendicant_bias):

\[\frac{t^{2}}{2}e^{t}\](Literally just entirely forgot what to do from here, one sec)

OpenStudy (sidsiddhartha):

now \[\int\limits_{0}^{t}F_1(\tau).F_2(t-\tau)d \tau\]

OpenStudy (mendicant_bias):

Alright, so just so we're on the same page, we're using two different instances of convolution? The first one to isolate F(s)/s, and the second one to find the laplace of F(s) itself in the form of F_19s) and F_2(s)?

OpenStudy (mendicant_bias):

\[\int\limits_{0}^{t}\frac{\tau^2}{2}e^{t-\tau}d \tau=e^t \int\limits_{0}^{t}\frac{\tau^2}{2}e^{\tau}d \tau\]

OpenStudy (sidsiddhartha):

yes just a typo that e^{- tau}

OpenStudy (mendicant_bias):

Evaluating the integral here using integration by parts,\[\int\limits_{0}^{t}\frac{\tau^2}{2}e^{-\tau}d \tau=-\frac{\tau^{2}e^{-\tau}}{2} -\int\limits_{0}^{t}\ -\tau e^{-\tau}\]

OpenStudy (mendicant_bias):

Integration by parts again,\[-\frac{\tau e^{-\tau}}{2}+\int\limits_{0}^{t}\tau e^{-\tau}d \tau=-\frac{\tau e^{-\tau}}{2}+\Bigg(-\tau e^{-\tau}-\int\limits_{0}^{t}-e^{-\tau}d \tau \Bigg)\]

OpenStudy (sidsiddhartha):

yes that looks fine :)

OpenStudy (mendicant_bias):

\[-\frac{\tau e^{- \tau}}{2}-\tau e^{-\tau}+e^{\tau}\]

OpenStudy (mendicant_bias):

Whoops, last exponent is still -tau

OpenStudy (mendicant_bias):

Evaluated (lol forgot to evaluate this whole time oh no)

OpenStudy (mendicant_bias):

\[e^{-t}\bigg[1-\frac{t}{2}-t\bigg]-\bigg[1-\frac{1}{2}\bigg]\]

OpenStudy (mendicant_bias):

Time to see how off this answer is from the real thing.

OpenStudy (mendicant_bias):

God, man, sometimes I feel like redoing this from scratch, real answer is \[e^{t}-\frac{1}{2}t^2-t-1\]Going to start over and try to finish this, then moving on to an integrodifferential equation problem, then dirac delta, then systems of linear diff eq, time is running out.

OpenStudy (sidsiddhartha):

man u are something :P